OpenStudy (crazysuperanya):

Medal for best answer Solve the system. y = x + 2 and y = -x + 0 (1, 3) (-1, 1) (-1, 3) (1, -1) no solution

OpenStudy (jsizzie101):

no solution

OpenStudy (anonymous):

plug in -1 in equation 1 into x

OpenStudy (anonymous):

(-1,1)

OpenStudy (anonymous):

Put the value of y from first equation into 2nd equation. So, x+2 = -x+0. Hence, x = -1. Put that back in 1st equation and you get y = -1 + 2 = 1. Answer is (-1,1).

OpenStudy (anonymous):

correct

OpenStudy (crazysuperanya):

Thank you

OpenStudy (anonymous):

you are welcome